La presentazione è in caricamento. Aspetta per favore

La presentazione è in caricamento. Aspetta per favore

STATO BARREL MU DT Roma 12/3/2007 F. Gasparini BARREL MU INSTALLATION IN YB0 UXC 7/3/2007.

Presentazioni simili

Presentazione sul tema: "STATO BARREL MU DT Roma 12/3/2007 F. Gasparini BARREL MU INSTALLATION IN YB0 UXC 7/3/2007."— Transcript della presentazione:

1 STATO BARREL MU DT Roma 12/3/2007 F. Gasparini BARREL MU INSTALLATION IN YB0 UXC 7/3/2007

2 YB+2 and +1 Installation in UX UX Chamber install. (8 ch/wheel) does not present particular problems The operations are easier than for the other sectors. Three wheels are in the cavern, complete of chambers and cables. All chambers were individually commissioned in surface before and after cabling The two negative wheels, are upstair, complete of chambers (only four per wheel the hor. sectors, out of eigth, and 2 in the feet of YB-2). 240/250 Installed All chambers passed the first individual Commissioning.

3 SX5 DT cabling of YB-1 is completed. Test of MiniCrates with final cabling was completed on March 21st. The entire YB-1 is under HV with the final cables and modules since March 23rd. YB-1 is now connected to the green barrack for final commissioning with cosmics. Data taking can start as soon as cooling is re-established to the racks, this week. Cabling of YB-2 might start in a week or so Massimo, Mimmo and the responsible for lowering in UX (Hubert Gerwig) agreed that chambers can be installed in stations 1 and 4 of the horizontal sectors without particular dangers or interference with the lifting gear. These chambers were installed from March 27th to 29th in YB-1 and YB-2. The chambers in YB-1 sector 1 and 7 will be connected to gas, cooling and cables in order to do a complete commissioning. The cables will be removed before lowering.

4 UX55 YB+2, YB+1 The connection of the HV cables from USC cannot start until the gas pipes are connected to the gas rack both for DT and RPCs. This operation requires welding the pipes in the cable chains to steel Pipes. Criotech, the firm in charge of this work, estimates that it will not be started before 4 to 5 weeks. (Priority given to YE Discs) We cannot put the optic fiber cables in the cable chains until this work is completed. Hence we cannot do any commissioning or test of YB+2 and YB+1 before the middle of May.

5 YB0 The test of MC in sectors 1 and 7 completed last week. YB0 is under gas since March 21st. The cable trays layout from YB0 to the balconies is still under discussion pending the definitions of the Tracker and ECAL services and cables routing. The most difficult area is on the first level. These cable trays power sectors 7, 8, 9, 12, 11 and 10. We do not know when they will be defined/installed The cable trays for HV to the chambers in sectors 1, 2, 3, 4, 5 and 6. are not so critical And could may be installed this week. The chambers in the top sectors might be put under HV by April 5th. The cable trays inside the YB0 trigger tunnel were installed the 27 of March. It should be possible to have the fiber optic cables installed by the end of this week. This would allow us to start setting up UX-USC communications.

6 PLANS & OBSTACLES A Wheel is a completely in dependent unit: it should be possible to operate, commission, synchronize each of them individually from the USC, just after the wheel is lowered. Actually this is impossible because the connection to USC is late No commissioning in UX before mid May. We are hardly negotiating an exception for YB0 : YB0 should pass the final commissioning BEFORE the cables and services for the inner detectors are laid down. They will prevent any intervention on the chambers and Minicrates. The lowering of the negative wheels was recently announced to be delayed at least until end of June. No time left for commissioning in the cavern. Commissioning will be done on surface per groups of three sectors,using the temporary Control Room set up for the Cosmic Challenge.

7 200420052006 33 82 204 BMu Installation Profile/Performance 0 50 100 150 200 250 Jun Aug Oct Dec FebAprJun Aug Oct Dec FebAprJun Aug Oct Dec FebAprJun Installed Chambers 240

8 SX5/UX5 HV interventions summary YB+2YB+1YB0YB-1YB-2Total HV interventions421029 Disconnected channels 121004 HV modules failures during pre-lowering test 4/120 A877(3 sent for repair to CAEN) total sent to CAEN 29/250 (12%) 1/12 A876 (sent for repair to CAEN)  we plan for next year (M&O B) for detector maintenance  special disconnect boxes for fixes with detector closed are in preparation (Jesus)  Chamber extraction tool ~ready for production (Massimo)

9 Cabling Update: YB0 YB0 cabling progressed quickly with two teams putting tremendous pressure on laying the fibers and testing the MC after cabling Thanks to the effort of F. Gonella, G. Masetti, M. Nervo, T. Rovelli and others, no delay was introduced in the YB0 cabling or UX schedule  Often, these details are lost in the official presentations


11 SX5, UXC Activities up to YB0 Commissioning Ready for checkout of YB0 commissioning setup Still to be ordered


13 Comments and Concerns UX and USC are not quite operational yet:  Racks are not connected to the cooling plant  The cable trays from YB0 towers to the balconies are still to be designed.  Fiber passage and protection in the UX-USC tunnels is still under design/discussion  Access to racks is difficult and safety must be improved DSS for UX and US racks not operational before end of May. We must setup a temporary system for smoke detection in UX racks and interlocks on cooling water flow/temperature

14 Summary YB0 data taking in UX is problematic since it requires full connection to USC, still uncharted territory. It should be completed before the ECAL/Tracker services are installed. This operation can start as soon as PP2 are installed (end of April?) Several items still to be installed in YB+2, YB+1 require access to the wheel with a scissor lift. This can be done when they are moved over the vactank but time will be very short. There is sufficient time for data taking with YB-1 from the green barrack but it is overlapping with YB0 data taking YB-1, YB-2 Lowering has been delayed > It can be done as soon as EB- installation is completed and after YB0 services on z- are installed (end of June).  We have some interesting months in front of us

15  ROS BOARDS (CIEMAT)  All (70) final modules produced, 21 now at CERN.  Next batch of ROS boards in March (all boards passing tests).  SECTOR COLLECTOR CRATES (CIEMAT)  10 final SC crates now at CERN.  tSC STATUS (Bologna)  First units available from February 15, full production completed by March.  Complete delivery at CERN (60 modules) end of April  DDU STATUS (Torino)  It is expected that DAQ bandwidth limit ( 200 MB/s) would be reached due to new proposed tSC data added on each event (4Bx).  move to 10 DDU  Planned production: 12 PCB, 5 mounted boards + 3 spares.  PCB production: delayed by 1-2 weeks. Now expected on 3 rd week of March.  Mounting: first 2 boards April 9/13. Rest will follow after tests.  Tests: 2-3 weeks. New tested boards by end of April. Full crate in May. STATUS OF “OFF DETECTOR” ELECTRONICS

16 DCS/DSS HW STATUS (Padova)  SLOW CONTROL MASTER Full production completed, 14 available, rest testing.  DCS SERVERS  Procured 5+1 Dell PCs. Ready for installation.  SC COLLECTOR CRATE VME CTR>  Full production (30 boards) finished. Under test.  Problem with Dell BIOS: possible solution move PCs to DAQ rack.  DSS  10 interlocks on A1517.  24 I/O to/from global DSS.  DCS interface based on Ethernet. Pending issues LV crate cooling tests: going Cable cutting lengths DCS/DSS USC cabling Most HW production seems to go on smoothly according to schedule. DDU 5 or 10 units?

17 CI SONO TRE EXTRACOSTI RECENTI : Fibre ottiche + DDU+ GAS ANALYSIS SYSTEM (CORE) + PROBLEMI DI HIRED MANPOWER DescrizioneMSF CORE MoU11,10 1o Contributo INFN extracosti2,20 2o Contributo INFN extracosti0,90 Contingenza (parte dei 800 kSF)0,00 Totale Accordato14,20 1. CORE Impegnato15,03 2. CORE da Impegnare0,04 3. Spares da sottrarre al CORE0,89 a.Totale speso a (1+2-3)14,18 4. Residuo Gare0,17 b.Totale Speso b (1+2-3-4)14,02 a.Residuo (accordato - speso)0,02 b.Residuo (Accordato - Speso)0,18 Sommario CORE - DT

18 CORE DARBO OCT 2006 Residuo assegnazione 2003 Residuo di assegnazione 2005 CORE DISPONIBILE 12 +28 + 40 KE = 120 KCHF 1. CORE Impegnato15,03 14,967 2. CORE da Impegnare0,04 0 3. Spares da sottrarre al CORE0,89 0,887 Totale speso a (1+2-3)14,18 14,020 Totale Accordato14,20

19 SCHEMA FINALE: Minicrate ===20~40 m=== patch panel ===~ 130m====== USCSlow Ctr Master COMMISSIONING: Minicrate ===20~40 m=== patch panel ===~ 20m ====== Test Stand Minicrate ===20~40 m=== patch panel ===~ 20m ====== ………………………………………………………. Le fibre sono provate e certificate dalla ditta che “connettorizza” secondo lo standard Internazionale. Si sono manifestate incertezze di comunicazione attribuite a polvere o a qualche Connettore difettoso ….pulire, cambiare conettore OK Nello MTCC sono in funzione 14 MC con lo schema nominale, 3 manifestano ancora incertezze e instabilita’, difficile attribuire l’origine: troppe variabili in gioco (rumore etc) Pero’ si decide di controllare a fondo le fibre prima dell’installazione in YB0. Il controllo di 100 fibre destinate all’installazione ne mostra 17 difettose: Perdita all’inserzione maggiore di 3~4 DBm Tutte le fibre sono fuori specifica : la pedita all’inserzione e’ di 2 DBM o piu’ invece di 0,67. Il controllo delle fibre lunghe nei bundles mostra le stesse peculiarita’. FIBRE OTT.

20 FIBRE 250 doppini da MC a Torri L. = 20 ~ 40 metri 10 bundles da 32 doppini da torri a USC L.= ~130 metri Gent. Dr. Bellato, le confermo che stiamo verificando tutte le alternative possibili, tra cui parlare con Lightec e Corning. La chiamo nel pomeriggio, per gli aggiornamenti Nel frattempo, come anticipato, Le inoltro il "report" del laboratorio: In relazione alla fornitura di cavi cablati e bretelle MTRJ ed al reclamo del cliente finale, abbiamo provveduto a controllare le 15 bretelle che sono ritornate presso il ns stabilimento,questa la situazione : le bretelle sono difettose, il difetto si manifesta con una rottura delle fibre o in superficie (assimilabile ad una frattura della fibra visibile al visual inspection) o con una rottura della fibra nei primi 3 mm all'interno della ferula MTRJ (visual inspection in superficie ok ma la fibra risulta interrotta) la rottura della fibra è provocata da un errato processo di resinatura (temperatura troppo alta), che ha prodotto un connettore fragile, in pratica dopo il nostro collaudo finale favorevole sia come visual inspection sia come insertion loss uno stress meccanico o termico può provocare con facilità la rottura della fibra. Etc etc…….. Cordiali saluti Silvia Picchi La ditta si dichiara pronta a rifare completamente il lavoro……… Sono state tutte riordinate da Corning che pero’ fornisce fibre E connettore


22 Expected DT ev.size = 7 Kbytes (140 Mbytes/s in each DDU). FIFOs can contain two events with 1 hit from each channels of one wheel. Bottle-neck in the DAQ (~ 200 Mbytes/s) DDU

23 Quando la DDU e' stata progettata non c'era ancora nell'architettura del sistema la possibilita‘ di trasferire dati dai Trigger sector collectors al DAQ (ROS e DDU). Venivano trasmessi alcuni bit di qualita’. cio’ non e’ sufficiente. La lezione del MTCC e’ stata infatti che per “debuggare” e tener sotto controllo il trigger l’informazione ottenibile via il DTTF non e’ adeguata, o sempre disponibile. E’ necessario trasferire via DAQ dati piu’ significativi. Cio’ implica un aumento dei dati forniti e trasmessi dalla DDU.

24 Expected DT ev.size = 7 Kbytes at 100 KHz => 700 Mbytes/sec ( 140 Mbytes/DDU) DDU is designed for 200 Mbytes/sec.(bottleneck of actual DAQ) In normal running addition of data from SC ( 576 bytes/event on +- 2 bx) at 100 KHZ gives 58 Mbytes/sec more, and saturate the 200 Mbytes. We should also remind that DAQ is considering to increase its bandwith from 200 to 300 Mbytes/sec.

25 We distinguish three types of data taking: 1) normal run : 4 Bxs => 384 bits/TSC => + 576 bytes to the DDU that saturates 2) initial run (2007/08 ) 20 Bxs => 1344 bits/TSC => + 2016 bytes to the DDU 3) set up: 128 BXS => 8320 bits/TSC =>+12480 bytes to the DDU So for initial running: 2 Kbytes  200 Mbytes from SC + 140 Mbytes data => 340 Mbytes/sec Affordable splitting SC + data on two DDU => 160 Mbytes/ DDU During set up we can reduce the L1A rate to fit our needs. Attualmente abbiamo 5 + 2 DDU,il loro raddoppio (10 DDU +1) ne richiede altre 4.

26 Gas analysis e Monitor della velocita’ di deriva Consente di analizzare contenuto di ossigeno e umidita all’ingresso e all’uscita di ognuna delle 250 camere. Il sistema e’ composto da cinque parti,una per ruota Iniettando il gas in cinque camere di riferimento si misura la velocita’ di deriva in assenza di campo magnetico ed in condizioni controllate. IL sistema e’ composto dai racks contenenti valvole e gauges + le VDC Entrambe le parti sono in assemblaggio/costruzione : il costo di entrambe e’ aumentato rispetto alle previsioni di parecchi anni fa.

27 The VDCs – part of the CMS gas system - sources: inlets to 5 wheels, outlet of 250 muon chambers, reference gas - compoundable to any of the (5+1) VDCs (serial + parallel) - piping network with 36 valves (hand-operated) VDC H2OH2O O2O2 US area UX area on wheel muon chambers mixer room.5.5 vent.5.5 n-way valve 50 lines to analyser 5.1.1 mixer room 50 5 1:3 p T 1

28 gas inlet 146 mm β-β- gas cover anode VDChambers Principle of operation and layout co drift time for 48 mm source: Sr 90 collimator Anode wire Sensitive region collimator events e-e- e-e- TDC stop start field shaping electrodes time CO=coincidence PM=photomultiplier 124 mm trigger (scintillator + 2 PMs)

29 HIRED MAN POWER: PERDIAMO A BREVE TERMINE TRE PERSONE Danilo Colonna ing. di Frascati attualmente PJS e braccio destro di Montecassiano Sandro di Vincenzo (CERN) : tecnico delle CMS infrastrutture (Herve’) lavora con noi da tre anni e viene richiamato al suo gruppo Jose’ Miguel: tecnico Spagnolo di CIEMAT permanente al CERN che rientra ( CIEMAT ha gia’ dato molto di piu’ del “dovuto”).

30 Gli extracosti sono stati divisi in modo che il costo di fibre e DDU sia tutto su INFN, Aumentando molto la percentuale Tedesca su “gas analysis + VDC e alterando un po’ le percentuali si suddivisione dei costi previsti per i cablaggi. A carico INFN sono : Fibre 60 ma alla fine 75 DDU 32 Cablaggi YB0 32 Sharing “gas analysis” 54 (oltre ai 56 gia’ assegnati) Hired Man power 100 totale 293 di cui 120 compresi nel CORE assegnato, 173 dalla contingenza

31 MISSIONI ESTERE E SBLOCCHI SJ PADOVA 55 / 261 21% BOLOGNA 63 / 288 22% TORINO 37 / 110 37% 2,2 mesi su 12 = 18 % SBLOCCHI BOLOGNA (28 SJ) chiede sblocco di 20 per cablaggi YB0 (Scambio con Aachen su finanziamento “gas analysis”, riduce extracosto INFN su “gas anlysis”) e di 2~3 per modifiche SC

Scaricare ppt "STATO BARREL MU DT Roma 12/3/2007 F. Gasparini BARREL MU INSTALLATION IN YB0 UXC 7/3/2007."

Presentazioni simili

Annunci Google